OpenStudy (anonymous):

Can someone check my answer please? On the graph of the equation 3x + 2y = 18, what is the value of the y-intercept? -9 -6 6 (My answer choice ) 9

OpenStudy (anonymous):

9

OpenStudy (solomonzelman):

1) set x=0 (i.e. substitute 0 instead of x) 2) solve for y OR 1) re-write the equation of the line in a y-intercept form. (i.e. y=mx+b) 2) set x=0 (i.e. substitute 0 instead of x)

OpenStudy (anonymous):

Could you explain please? I used x=2 and plugged it in for x and got 6.

OpenStudy (solomonzelman):

you have to use x=0.I will explain why....

OpenStudy (solomonzelman):

y-intercept is when the graph of the line hits the y-axis, and that would happen when x=0 (because y-axis and the line x=0 is the same thing) do you agree?

OpenStudy (solomonzelman):

(not necessarily the graph of a LINE, but in this case, a line)

OpenStudy (anonymous):

Oh okay! I understand thank you :) Would you mind helping me with another? I'm so terrible at math :(

OpenStudy (solomonzelman):

sure, if I get time

OpenStudy (anonymous):

blah blah XD

OpenStudy (anonymous):

Money mob guy, lol if I post it on another post and you help I'll give you a metal :)